OpenStudy (blackops2_143):

What is -1 2/3 divided by 5/6? A. -1 7/18 B.-2 C. 2 D. -1 1/6

OpenStudy (blackops2_143):

I will medal and fan any help.

OpenStudy (kendricklamar2014):

-1 2/3 = -5/3 \[\frac{ -5 }{ 3 } \div\frac{ 5 }{ 6 } = -2\]

OpenStudy (ibrafacts):

yup its -2